# EA-KD: Entropy-based Adaptive Knowledge Distillation

### v2

@@ Abstract
Knowledge distillation (KD) enables a smaller “student” model to mimic a larger “teacher” model by transferring knowledge from the teacher’s output or features. However, most KD methods treat all samples uniformly, overlooking the varying learning value of each sample and thereby limiting effectiveness. In this paper, we propose Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that prioritizes learning from valuable samples. EA-KD quantifies each sample’s learning value by strategically combining the entropy of the teacher and student output, then dynamically reweights the distillation loss to place greater emphasis on high-value samples. Extensive experiments across diverse KD frameworks and tasks–including image classification, object detection, and large language model (LLM) distillation–demonstrate that EA-KD consistently enhances performance, achieving state-of-the-art results with negligible computational cost. Our code will be publicly available.
11footnotetext: Equal contribution22footnotetext: Corresponding author: camhero@gmail.com
@@ 1 Introduction
The increasing size of state-of-the-art (SOTA) deep learning models in computer vision (CV) [32, 21, 6], and large language models (LLMs) [5, 25] pose challenges for deployment in resource-constrained settings. Knowledge Distillation (KD) [9] offers a solution by training a smaller “student” model to mimic a larger “teacher” model, using both ground-truth labels and the teacher’s “knowledge” (e.g. softened logits or feature representations) to achieve similar performance in a compact form.
@@
Figure 1: Training Time per Batch (ms) vs. Testing Accuracy (%) on CIFAR-100. Highlighting EA methods’ enhanced performance with minimal additional training cost.
@@
SOTA KD methods have explored diverse knowledge forms and structural modifications to refine knowledge transfer [9, 26, 1, 24, 38, 11, 19]. However, most methods distill uniformly across all samples, operating under the assumption that each sample has equal importance and overlooking their varying learning needs. Recent advancements, such as Instance-T [16], have shown that assigning unique temperatures to each sample outperforms the uniform temperature method Global-T. This highlighted the benefits of adapting to each sample’s distinct characteristics. Building on this, we hypothesize that emphasizing samples rich in valuable knowledge [^1] Throughout this paper, we consider samples with high entropy as containing valuable knowledge. can further optimize the distillation process. This mimics how human students learn better when key points are highlighted by teachers, and thus, this targeted approach allows the student model to achieve better comprehension by focusing on more informative samples. As we will demonstrate, the uniform distillation in most KD methods also tends to overshadow these important samples, limiting the efficiency of knowledge transfer.
@@
The question then arises: How can we identify the most valuable samples for learning? Entropy, the core concept in information theory that quantifies the uncertainty or information of a random variable [28], may be well-suited for this role. While a prior multi-teacher KD study utilized inverse entropy to measure the reliability of teachers’ output [13], this method does not apply to the more prevalent single-teacher setting. Moreover, relying solely on the teacher’s entropy restricts the adaptability to the evolving learning dynamics of the student, leaving the full potential of entropy in KD under-explored. Therefore, we propose leveraging entropy to quantify the learning value of each sample in KD, as high-entropy outputs shall correspond to greater information contents that are crucial for learning. Our preliminary analysis reveals that higher entropy samples [^2] Results for high teacher entropy samples are shown here; high student entropy plots are in the Sec. A.1. (i) correlate with larger teacher-student accuracy gaps (Fig. 2(a)) and (ii) often lie near class boundaries in t-SNE visualizations [33] (Fig. 2(b)), suggesting these informative samples not only offer valuable learning opportunities but are also pivotal in defining decision boundaries. Hence, entropy can serve as a reliable metric for identifying valuable samples that are crucial for learning in KD.
@@
Figure 2: High-Entropy Samples in KD. (a) Higher teacher entropy correlates with larger accuracy gaps in KD, while EA-KD maintains closer alignment. (b) The top 10% of high entropy samples (denoted by triangles) cluster near decision boundaries, representing critical knowledge essential for classification.
@@
Leveraging this insight, adapting the focus of KD to valuable samples through entropy reweighting should fill the need for enhanced distillation. We conducted a preliminary experiment that compared the performance of reweighting with teacher ($H^{\mathcal{T}}$) and student ($H^{\mathcal{S}}$) entropy, along with their linear-inverted variants ($H_{\text{ub}}-w$), where high-entropy samples received lower weight. Here, $H_{\text{ub}}$ is the upper bound of entropy ($log100$ for CIFAR-100). As shown in Tab. 1, reweighting with either $H^{\mathcal{T}}$ or $H^{\mathcal{S}}$ significantly improved KD accuracy, while inverted reweighting led to decreased performance. This supports our hypothesis that focusing on samples with valuable knowledge enhances student learning, with $H^{\mathcal{T}}$ proving more effective due to the teacher’s more reliable assessment of sample value.
@@
Figure 3: Box Plot of Student vs. Teacher Entropy for Top 10% High Entropy Samples Across Epochs. $H^{\mathcal{S}}$ shows significant variations for high $H^{\mathcal{T}}$ samples, highlighting the student’s different views from the teacher. EA-KD achieves improved mimicry of the teacher with more stable entropy and better alignment.
In light of the above analysis, we introduce Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that directs learning toward high-value samples by reweighting the distillation loss. Leveraging both teacher and student entropies, EA-KD emphasizes samples that encapsulate critical knowledge as identified by the teacher, while dynamically adapting to the student’s evolving needs. Intuitively, this method addresses a fundamental limitation in standard KD frameworks, which often overlook the unique learning value of each sample and tend to bias towards simpler knowledge ones. Furthermore, EA-KD can be seamlessly integrated into most KD frameworks, enhancing their performance with negligible computational cost (Fig. 1). Extensive experiments on image classification, object detection, and LLM distillation demonstrate its efficacy and versatility across diverse KD frameworks. Our main contributions are summarized as follows.
  • We propose EA-KD, a plug-and-play KD method that adaptively reweights the distillation loss to prioritize samples valuable for learning, enabling a more nuanced and efficient knowledge transfer process.
  • We show that high-entropy samples carry critical knowledge in KD, and propose an entropy-based reweighting factor that integrates both teacher and student entropy for a tailored and dynamic focus on these samples.
  • We show that EA-KD consistently enhances performance in both logit- and feature-based KD across CV and LLM tasks, achieving SOTA results with negligible cost.
@@ 2 Related Work
Logit and Feature Distillation. Logit distillation [9, 38, 11] aligns the softened output logits of the teacher and student, valued for its simplicity and broad applicability. On the other hand, feature distillation minimizes divergence in intermediate feature representations, offering enhanced learning but often with higher computational costs [26, 1, 19]. Both pathways have achieved SOTA performance across tasks and domains. However, most of them typically adopt a static distillation scheme, such as treating all samples uniformly. Adaptive distillation addresses this limitation by introducing more dynamic knowledge transfer processes [30, 16, 20, 22, 13].
Adaptive Distillation. These methods refine knowledge transfer by adapting to data characteristics, student capabilities, or training dynamics. For instance, LS [30] uses Z-score standardization on logits, allowing the student to focus on essential relationships in the teacher’s output. CTKD [16] and MKD [20] adjust the temperature parameter to meet student needs, with Instance-T [16] highlighting the effectiveness of sample-wise adjustments. RW-KD [22] improves distillation by using meta-learning to find optimal weights for each sample. However, these methods often increase complexity and computational costs [16, 20, 22] or are limited to a specific type of KD [30]. In contrast, EA-KD offers a straightforward, more interpretable approach that is versatile across diverse KD frameworks.
Entropy in KD. Previous work introduced an entropy-based metric to quantify knowledge in KD, measuring how much visual information is retained or discarded [2]. Inspired by this, we adopted the entropy of model output to identify valuable samples in KD. AKD [13], which also leverages entropy in KD, adjusts weights based on teacher entropy in multi-teacher distillation, placing greater emphasis on low-entropy predictions. On the contrary, our approach prioritizes high-entropy samples in single-teacher distillation. While low entropy may indicate reliable knowledge in multi-teacher schemes, we view lower entropy as simpler knowledge and is less beneficial for student learning in a single-teacher context, as demonstrated in our preliminary analysis and experiments in Sec. 1.
@@ 3.1 Preliminaries
$H_{n}=-\sum_{j=1}^{C}\sigma(z_{n,j})\log(\sigma(z_{n,j}))\,,$ (1)
where $\sigma(\cdot)$ denotes the softmax function, $z_{n,j}$ is the logit for class $j$ of sample $x_{n}$, and $C$ is the number of classes.
@@
$p_{i}(T)=\sigma(z,T)_{i}=\frac{\exp(\frac{z_{i}}{T})}{\sum_{j=1}^{C}\exp(\frac{z_{j}}{T})}\,,$ (2)
@@
$\begin{split}L_{\text{KD}}&=\text{KL}(p^{\mathcal{T}}(T)\|p^{\mathcal{S}}(T))\cdot T^{2}\\ &=\sum_{j=1}^{C}p^{\mathcal{T}}_{j}(T)\log\left(\frac{p^{\mathcal{T}}_{j}(T)}{p^{\mathcal{S}}_{j}(T)}\right)\cdot T^{2}\,,\end{split}$ (3)
@@
$\begin{split}L_{\text{KD}}=\sum_{j=1}^{C}p^{\mathcal{T}}_{j}\log\left(\frac{p^{\mathcal{T}}_{j}}{p^{\mathcal{S}}_{j}}\right).\end{split}$ (4)
Figure 4: Illustration of the Uniform Distillation Scheme in Standard KD (Left) and Entropy-based Adaptive Reweighting in EA-KD (Right). The uniform distillation scheme overlooks the learning value of individual samples, while EA-KD prioritizes valuable samples based on both teacher and student assessments.
@@ 3.2 EA-KD
Limitations in Standard KDs. Most logit- and feature-based KD methods typically apply a uniform distillation across all samples, overlooking their unique learning value. This oversight can cause the model to over-prioritize simpler samples at the expense of more valuable, high-entropy ones. Taking KLD—the main loss function for logit-based methods—as an example, consider a student initialized with a uniform distribution $p^{\mathcal{S}}$ where $p^{\mathcal{S}}_{i}=\frac{1}{C}\,\forall\,i$. For a low-entropy sample with a teacher output $p^{\mathcal{T}}_{\text{low}}$ where $p^{\mathcal{T}}_{\text{low},j}\approx 1$ and $p^{\mathcal{T}}_{\text{low},i}\approx 0\,(i\neq j)$, the KLD simplifies to:
$\begin{split}\text{KL}(p^{\mathcal{T}}_{\text{low}}\|p^{\mathcal{S}})&\approx p^{\mathcal{T}}_{\text{low},j}\cdot\log\left(\frac{p^{\mathcal{T}}_{\text{low},j}}{p^{\mathcal{S}}_{j}}\right)\\ &=\log\left(\frac{1}{p^{\mathcal{S}}_{j}}\right)\\ &=\log(C).\end{split}$ (5)
@@
$\begin{split}\text{KLD}(p^{\mathcal{T}}_{\text{high}}\|p^{\mathcal{S}})&\approx\sum_{j=1}^{C}\frac{1}{C}\cdot\log\left(\frac{1/C}{1/C}\right)\\ &=0.\end{split}$ (6)
@@
This inequality highlights KLD’s overemphasis on low-entropy samples, causing them to dominate the loss over high-entropy ones. Similarly, MSE loss in feature-based methods also biases learning toward high-magnitude activations, which typically correspond to low-entropy samples. Such imbalance shifts the learning focus toward simpler samples, hindering the transfer of knowledge crucial for learning, especially during early training stages (refer to the overshadowing effect in the left part of Fig. 4).
Entropy-Based Reweighting. As discussed in Sec. 1, entropy can serve as a measure of the sample learning value in KD. In practice, we soften the entropy with an alternative temperature $T^{\prime}$ to better capture the value, expressed as:
$H_{n}=-\sum_{j=1}^{C}p_{n,j}(T^{\prime})\log(p_{n,j}(T^{\prime}))\,,$ (8)
where $p_{n,j}(T^{\prime})$ is the temperature-scaled probability of class $j$ for sample $x_{n}$.
To effectively target and emphasize samples valuable for the student, EA-KD’s reweighting factor $w_{\text{EA}}$ is formulated using two components: a base term $w_{\text{base}}$ and an interaction term $w_{\text{interact}}$. The base term captures the inherent value of each sample based on the teacher’s output entropy, defined as:
@@
where $H^{\mathcal{T}}_{n}$ is the entropy of the teacher’s prediction for sample $x_{n}$, and $H_{\text{ub}}=\log(C)$ denotes the upper bound of entropy for $C$ classes. The interaction term $w_{\text{interact}}$, on the other hand, captures the interplay between the teacher and student perspectives by taking the normalized product of their entropies:
@@
This reformulation expresses $w_{\text{EA},n}$ as a product of the $H^{\mathcal{T}}_{n}$ and a scaling factor that depends on the $H^{\mathcal{S}}_{n}/H_{\text{ub}}$. In this way, the teacher’s assessment of sample value is adaptively adjusted based on the student’s perspective. Depending on the entropy values, $w_{\text{EA}}$ behaves as follows:
@@
$w_{\text{EA},n}=$ 00 0 if $H^{\mathcal{T}}_{n}\rightarrow 0\quad\forall H^{\mathcal{S}}_{n}$ (13c)
@@
Table 2: Results on CIFAR-100. Accuracy (%) of EA-methods vs. originals across teacher-student pairs, with relative improvements highlighted. Avg. $\Delta$ shows average improvement across methods and model pairs. Best results are bolded, and second-best are underlined.
Type Teacher ResNet32$\times$4 WRN-28-4 WRN-40-2 VGG13 VGG13 ResNet50 ResNet32$\times$4
@@
$\begin{split}L_{\text{EA-KD}}&=\sum_{n=1}^{N}w_{\text{EA},n}\cdot L_{\text{KD},n}\\ &=\sum_{n=1}^{N}w_{\text{EA},n}\cdot\text{KL}(p^{\mathcal{T}}_{n}\|p^{\mathcal{S}}_{n})\,,\end{split}$ (14)
where $L_{\text{KD},n}$ is the distillation loss for sample $x_{n}$, and $p^{\mathcal{T}}_{n}$ and $p^{\mathcal{S}}_{n}$ denote the teacher and student’s softened probability for $x_{n}$, respectively. Consequently, EA-KD introduces a more nuanced and dynamic knowledge transfer to standard KD frameworks, ensuring that valuable samples receive enhanced focus throughout training.

### v3

@@ Abstract
Knowledge distillation (KD) enables a smaller “student” model to mimic a larger “teacher” model by transferring knowledge from the teacher’s output or features. However, most KD methods treat all samples uniformly, overlooking the varying learning value of each sample and thereby limiting effectiveness. In this paper, we propose Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that prioritizes learning from valuable samples. EA-KD quantifies each sample’s learning value by strategically combining the entropy of the teacher and student output, then dynamically reweights the distillation loss to place greater emphasis on high-entropy samples. Extensive experiments across diverse KD frameworks and tasks–including image classification, object detection, and large language model (LLM) distillation–demonstrate that EA-KD consistently enhances performance, achieving state-of-the-art results with negligible computational cost. Our code is available at https://github.com/cpsu00/EA-KD.
11footnotetext: Equal contribution.22footnotetext: Bin Pu and Shin-Jye Lee are corresponding authors (eebinpu@ust.hk, camhero@gmail.com).
@@ 1 Introduction
The growing size of state-of-the-art (SOTA) deep learning models poses challenges for deployment in resource-constrained settings. Knowledge Distillation (KD) [11] offers a solution by training a smaller “student” model to mimic a larger “teacher” model, using both ground-truth and the teacher’s “knowledge” (e.g. logits or feature representations) to achieve similar performance in a compact form. KD has been widely applied across domains, including computer vision (CV) [40, 1, 31], large language models (LLMs) [35, 13, 9, 15], and medical imaging [21, 22, 23, 8].
@@
Figure 1: Training Time per Batch (ms) vs. Accuracy (%) of EA-Methods and Baselines on CIFAR-100.
@@
Advanced KD methods have explored diverse knowledge forms and structural modifications to refine knowledge transfer [11, 33, 2, 30, 47, 14, 26]. However, most methods distill uniformly across all samples, operating under the assumption that each sample has equal importance and overlooking their varying learning value. Recent advancements, such as Instance-T [20], have shown that assigning unique temperatures to each sample outperforms the uniform temperature method Global-T. This highlighted the benefits of adapting to each sample’s distinct characteristics. Building on this, we hypothesize that emphasizing samples rich in valuable knowledge [^1] Throughout this paper, we consider samples with high entropy as containing valuable knowledge. can further optimize the distillation process. This mirrors how human students learn better when key points are highlighted by teachers. Such a strategy enables the student model to focus on more informative samples, leading to improved performance. However, as we will show, the uniform distillation scheme in most KD methods often overlooks these critical samples, thereby limiting the efficiency of knowledge transfer.
@@
The question then arises: How can we identify the most valuable samples for learning? Entropy, the core concept in information theory that quantifies the uncertainty or information of a random variable [36], may be well-suited for this role. Prior methods have incorporated entropy in KD to adjust weighting [17] or refine logit predictions [49]; however, these approaches are typically restricted to specific scenarios (e.g., multi-teacher or logit-based KD) and primarily rely on teacher entropy, leaving the broader potential of entropy in KD underexplored. Therefore, we propose leveraging entropy to quantify the learning value of each sample in KD, as high-entropy outputs shall correspond to greater information contents that are crucial for learning. Our preliminary analysis reveals that higher entropy samples [^2] Results for high teacher entropy samples are shown here; high student entropy plots are in the Sec. B.1. (i) correlate with larger teacher-student accuracy gaps (Fig. 2(a)) and (ii) often lie near class boundaries in t-SNE visualizations [41] (Fig. 2(b)), suggesting these informative samples not only offer valuable learning opportunities but are also pivotal in defining decision boundaries. Thus, entropy can serve as a reliable metric for identifying the most valuable samples in KD.
@@
Figure 2: High-Entropy Samples in KD. (a) Higher teacher entropy correlates with larger accuracy gaps in KD, while EA-KD maintains closer alignment. (b) The top 10% high entropy samples (denoted by triangles) cluster near decision boundaries, representing critical knowledge essential for classification.
@@
Leveraging this insight, adapting the focus of KD to valuable samples should fill the need for enhanced distillation. To validate this, we conducted a preliminary experiment that compared the performance of reweighting with teacher ($H^{\mathcal{T}}$) and student ($H^{\mathcal{S}}$) entropy, along with their linear-inverted variants ($H_{\text{ub}}-w$), where high-entropy samples received lower weight. Here, $H_{\text{ub}}$ denotes the upper bound of entropy (i.e., $\log 100$ for CIFAR-100). As shown in Tab. 1, reweighting with either $H^{\mathcal{T}}$ or $H^{\mathcal{S}}$ significantly improved KD accuracy, while inverted reweighting led to decreased performance. This supports our hypothesis that focusing on samples with valuable knowledge enhances student learning, with $H^{\mathcal{T}}$ proving more effective due to the teacher’s more reliable assessment of sample value.
@@
Figure 3: Box Plot of $H^{\mathcal{S}}$ vs. $H^{\mathcal{T}}$ for Top 10% High $H^{\mathcal{T}}$ Samples. $H^{\mathcal{S}}$ shows notable variation, reflecting the student’s learning dynamics and differing perspective from the teacher. EA-KD enhances mimicry with more stable and aligned entropy.
In light of the above analysis, we introduce Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that enhances distillation by reweighting the loss toward high-value samples. Leveraging both teacher and student entropies, EA-KD emphasizes samples that encapsulate critical knowledge as identified by the teacher, while dynamically adapting to the student’s evolving needs. This method directly addresses a fundamental limitation in standard KD frameworks, which often overlook the unique learning value of each sample and tend to bias towards simpler knowledge. Furthermore, EA-KD can be seamlessly integrated into most KD frameworks, enhancing their performance with negligible computational cost (Fig. 1). Extensive experiments on image classification, object detection, and LLM distillation demonstrate its efficacy and versatility across diverse KD frameworks. Our main contributions are summarized as follows:
  • We reveal that high-entropy samples carry critical knowledge in KD and propose an entropy-based reweighting factor that integrates both teacher and student entropy to provide a dynamic and tailored learning focus.
  • We introduce EA-KD, a plug-and-play KD method that adaptively reweights the distillation loss to prioritize valuable samples, enabling more effective and efficient knowledge transfer.
  • We demonstrate that EA-KD consistently improves performance across logit- and feature-based KD methods, achieving SOTA results on both CV and LLM tasks with minimal computational overhead.
@@ 2 Related Work
Logit and Feature Distillation. Logit distillation [11, 47, 14] aligns the softened output logits of the teacher and student, valued for its simplicity and broad applicability. On the other hand, feature distillation minimizes divergence in intermediate feature representations, offering enhanced learning but often with higher computational costs [33, 2, 26]. Both pathways have achieved SOTA performance across tasks and domains. However, most of them typically adopt a static distillation scheme, such as treating all samples uniformly. Adaptive distillation addresses this limitation by introducing more dynamic knowledge transfer processes [28, 46, 20, 38, 17, 49, 48]
Adaptive Distillation. These methods improve knowledge transfer by dynamically adjusting knowledge at different levels. For sample-level, RW-KD [28] employs meta-learning to optimize weights for each sample, which introduces high computational overhead. PAD [46] showed that traditional hard-mining weighting is unsuitable for KD and instead prioritizes samples with low uncertainty and small teacher-student gaps. This shares similarities with EA-KD, as we will show that high-entropy samples often exhibit lower KLD. However, PAD relies on an auxiliary estimation branch, whereas EA-KD offers a more interpretable and efficient approach by directly utilizing entropy. For logit-level, CTKD [20] and LS [38] dynamically adjust the temperature parameter $T$ to refine knowledge transfer, but remain limited to logit-based KD. Instead, EA-KD’s sample-wise reweighting ensures broader compatibility across KD frameworks. Importantly, EA-KD and these methods serve distinct yet complementary roles, combining them could further improve performance, as we will show.
Entropy in KD. Cheng et al. introduced an entropy-based metric to quantify knowledge retention in KD [3]. Inspired by this, we utilize entropy to identify valuable samples in KD. AKD [17] assign higher weights to low-entropy teacher predictions in multi-teacher settings. However, such weighting can degrade performance in the more common single-teacher settings (Tab. 1). DynamicKD [49] refines logit-level knowledge through entropy correction similar to CTKD and LS, and is also constrained to logit-based KD. TTM [48] removes the student temperature, revealing an inherent Rényi entropy regularization, while WTTM further emphasizes the uncertain samples. In contrast, EA-KD actively leverages both teacher and student entropy for dynamic weighting, yielding stronger performance [^3] See Sec. A.1 for a more detailed comparison..
@@ 3.1 Preliminaries
$H_{n}=-\sum_{i=1}^{C}\sigma(z_{n,i})\log(\sigma(z_{n,i}))\,$, (1)
where $\sigma(\cdot)$ denotes the softmax function, $z_{n,i}$ is the logit for class $i$ of sample $x_{n}$, and $C$ is the number of classes.
@@
$p_{i}(T)=\sigma(z,T)_{i}=\frac{\exp(\frac{z_{i}}{T})}{\sum_{k=1}^{C}\exp(\frac{z_{k}}{T})}\,$, (2)
@@
[…]
@@
$\begin{split}L_{\text{KD}}=\sum_{i=1}^{C}p^{\mathcal{T}}_{i}\log\left(\frac{p^{\mathcal{T}}_{i}}{p^{\mathcal{S}}_{i}}\right).\end{split}$ (4)
Figure 4: Illustration of the Uniform Distillation Scheme in Standard KD (Left) and Entropy-based Reweighting in EA-KD (Right). Standard KD not only overlook the varying learning value of individual samples but also bias toward low-entropy (easier) ones. In contrast, EA-KD effectively guides learning toward valuable samples based on both teacher and student assessments.
@@ 3.2 EA-KD
Limitations in Standard KDs. Most logit- and feature-based KD methods [11, 47, 14, 2, 26, 33] treat all samples uniformly, overlooking their unique learning value. This oversight can cause the model to over-prioritize simpler samples at the expense of more valuable, high-entropy ones. Taking KLD—the main loss function for logit-based methods—as an example, consider a student initialized with a uniform distribution $p^{\mathcal{S}}$ where $p^{\mathcal{S}}_{i}=\frac{1}{C}\,\forall\,i$. For a low-entropy sample with a teacher output $p^{\mathcal{T}}_{\text{low}}$ where $p^{\mathcal{T}}_{\text{low},j}\approx 1$ and $p^{\mathcal{T}}_{\text{low},i}\approx 0\,(i\neq j)$, the KLD becomes:
$\begin{split}\text{KLD}(p^{\mathcal{T}}_{\text{low}}\|p^{\mathcal{S}})&\approx p^{\mathcal{T}}_{\text{low},j}\cdot\log\left(\frac{p^{\mathcal{T}}_{\text{low},j}}{p^{\mathcal{S}}_{j}}\right)\\ &=\log\left(\frac{1}{p^{\mathcal{S}}_{j}}\right)\\ &=\log(C).\end{split}$ (5)
@@
$\begin{split}\text{KLD}(p^{\mathcal{T}}_{\text{high}}\|p^{\mathcal{S}})&\approx\sum_{i=1}^{C}\frac{1}{C}\cdot\log\left(\frac{1/C}{1/C}\right)\\ &=0.\end{split}$ (6)
@@
This inequality highlights KLD’s inherent bias toward low-entropy samples, which dominate the loss over the valuable, high-entropy ones. Similarly, the MSE loss in feature-based KD also biases learning toward high-magnitude activations, which typically correspond to low-entropy samples. Such imbalance shifts the learning focus toward simpler samples, hindering the transfer of knowledge crucial for learning, especially during early training stages (refer to the overshadowing effect in the left part of Fig. 4).
Entropy-based Reweighting. As discussed in Sec. 1, entropy can serve as a measure of the sample learning value in KD. In practice, we soften the entropy with an alternative temperature $T^{\prime}$ to better reflect sample value:
$H_{n}=-\sum_{i=1}^{C}p_{n,i}(T^{\prime})\log(p_{n,i}(T^{\prime}))\,$, (8)
where $p_{n,i}(T^{\prime})$ is the temperature-scaled probability of class $i$ for sample $x_{n}$.
To dynamically emphasize valuable samples, EA-KD’s reweighting factor $w_{\text{EA}}$ is formulated using two components: a base term $w_{\text{base}}$ and an interaction term $w_{\text{interact}}$. The base term captures the inherent value of each sample based on the teacher’s output entropy, defined as:
@@
where $H^{\mathcal{T}}_{n}$ denotes the entropy of the teacher’s prediction for sample $x_{n}$, and $H_{\text{ub}}=\log(C)$ is the upper bound of entropy for $C$ classes. The interaction term $w_{\text{interact}}$, on the other hand, captures the interplay between the teacher and student perspectives by taking the normalized product of their entropies:
@@
This reformulation expresses $w_{\text{EA},n}$ as the product of $H^{\mathcal{T}}_{n}$ and a scaling factor that depends on the $H^{\mathcal{S}}_{n}$. Thus, $w_{\text{EA},n}$ can be regarded as the teacher’s assessment of sample value, adaptively adjusted based on the student’s perspective. Depending on the entropy values, $w_{\text{EA}}$ behaves as follows:
@@
$w_{\text{EA},n}=$ $0$ if $H^{\mathcal{T}}_{n}\rightarrow 0\quad\forall H^{\mathcal{S}}_{n}$ (13c)
@@
Table 2: Results on CIFAR-100. Accuracy (%) of EA-methods vs. baselines across teacher-student pairs, with relative improvements highlighted. Avg. $\Delta$ shows average improvement across methods and model pairs. Best results are bolded, and second-best are underlined.
[…]
@@
$\begin{split}L_{\text{EA-KD}}&=\sum_{n=1}^{N}w_{\text{EA},n}\cdot L_{\text{KD},n},\end{split}$ (14)
where $L_{\text{KD},n}$ is the distillation loss for sample $x_{n}$. As a result, EA-KD enhances standard KD methods by facilitating a more nuanced and adaptive transfer of knowledge, ensuring that informative samples receive increased focus throughout training.